Ct =a+bYt-1...(1)

Kt=CYt-1...(2)

It=Kt - Kt-1= Kt - CYt-2...(3)

Gt =G....(4)

Yt = Ct + It + Gt

t In terms of Yt-1, Yt-2, and the parameter of the model.


1
Expert's answer
2021-03-31T07:28:19-0400
"solution"

"c_t-a=b[Y_{t-1}]"


"K_t=C[Y_{t-1}]"


"L_t-K_t=C[Y_{t-2}]"


"G_t=G"


"Y_{t+1 }= G_{t^2}"